# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 5.6 Fundamental Tradeoffs
Four additional trade-offs pervade reinforcement learning. First, exploration versus exploitation: should the agent act on its current best estimate or gather information to improve future decisions? Lai and Robbins (1985) establish the fundamental lower bound: any consistent policy must incur regret at least logarithmic in the number of periods. Naive exploration ($\varepsilon$-greedy) requires samples exponential in the horizon; strategic exploration (UCB, optimism in the face of uncertainty) reduces this to polynomial (Auer et al., 2002a), formalizing the value of targeted experimentation. [^99] The exploration-exploitation tradeoff is the subject of the bandits chapter, where the multi-armed bandit framework provides the sharpest analysis. In full MDPs, exploration is harder because the agent must learn not just reward distributions but also transition dynamics, compounding the information requirement.
Second, model-based versus model-free: model-based methods learn a transition model $\hat{P}(s^{\prime}|s,a)$ and plan with it (Sutton, 1990); model-free methods learn value functions or policies directly from transitions. The Dyna architecture (Sutton, 1990) bridges these by generating simulated experience from the learned model to supplement real transitions. Model-based methods are sample-efficient (each transition updates the entire model, which improves value estimates for all states) but suffer asymptotic bias if the model class is misspecified; model-free methods are asymptotically unbiased but sample-inefficient, using each transition for a single gradient step.100100100Model misspecification in RL is the analog of omitted variable bias in econometrics: if the learned model omits relevant state variables or misspecifies the functional form of transitions, the resulting policy is biased regardless of sample size. Moerland et al. (2023) provide a comprehensive survey of model-based RL, analyzing the model-bias versus sample-efficiency tradeoff across method families.
@@ 6.1 The Moving Target Problem
Fujimoto et al. (2022) formalize this observation. Let $Q^{\pi}$ denote the true value function for policy $\pi$, let $\Delta(s,a)=Q(s,a)-Q^{\pi}(s,a)$ denote the value error, and let $\varepsilon(s,a)=Q(s,a)-(r+\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[Q(s^{\prime},a^{\prime})])$ denote the Bellman error. Substituting the definition of $Q^{\pi}$ yields the key identity: $\varepsilon(s,a)=\Delta(s,a)-\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[\Delta(s^{\prime},a^{\prime})]$. The Bellman error is a difference of value errors at consecutive states, not the value error itself. If the errors $\Delta(s,a)$ and $\Delta(s^{\prime},a^{\prime})$ are correlated across time—the network is wrong in the same direction at successive state-action pairs—they cancel in the difference and the Bellman error is small regardless of how large the individual errors are.101101101In the extreme case, shifting all Q-values by the constant $c/(1-\gamma)$ leaves the Bellman error unchanged at zero while increasing value error by $c/(1-\gamma)$.
The second failure mode is specific to finite datasets: Fujimoto et al. (2022, Corollary 1) show that over an incomplete dataset, zero Bellman error is consistent with arbitrarily large value error, because the network can fit unobserved successor pairs to whatever values make the observed residuals vanish.102102102The Bellman equation uniquely identifies $Q^{\pi}$ when enforced over the entire MDP, but over an incomplete dataset it admits infinitely many solutions. Whenever a transition $(s^{\prime},a^{\prime})$ that is reachable from the dataset is not itself in the dataset, the network is free to set $Q(s^{\prime},a^{\prime})$ at unobserved pairs to whatever value makes the residual vanish on the observed ones, unconstrained by any loss. A network can thus reach near-zero training loss while the value function remains arbitrarily inaccurate over the full state space.
The dual failure mode appears in policy gradient methods: Ilyas et al. (2020) find that even when PPO’s surrogate objective improves monotonically, episode return can plateau or decline, because the surrogate gradient is poorly aligned with the gradient of the true return.103103103Ilyas et al. (2020) examine the surrogate objective used in Proximal Policy Optimization (Schulman et al., 2017). The PPO clipping mechanism is designed to keep policy updates within a trust region by bounding the probability ratio $\pi_{\theta}(a|s)/\pi_{\theta_{\text{old}}}(a|s)$. The gradient of the surrogate is poorly aligned with the gradient of the true return, particularly in later training phases where the policy has diverged from the behavior policy used to collect the replay data. The loss metric that practitioners monitor throughout training is measuring something other than what they care about.
@@ 6.2 The Reproducibility Crisis and Sensitivity to Random Seeds
Henderson et al. (2018) trained five leading policy gradient algorithms (PPO, TRPO, DDPG, TD3, SAC) on six MuJoCo benchmark environments, holding all hyperparameters fixed and varying only the random seed. The resulting learning curves from different seeds were non-overlapping: a seed that performed well under one algorithm performed comparably to a different algorithm’s best seeds, making cross-algorithm comparison unreliable.104104104Differences as large as 2,000 points in final episode return arose from seed variation alone. Agarwal et al. (2021b) quantify the damage: comparing point estimates from 5 runs per task on Atari 100k yields Type I error exceeding 50%, meaning a random noise injection appears beneficial in half of all comparisons.
Agarwal et al. (2021b) propose the interquartile mean (IQM) as a replacement for mean and median when comparing algorithms. The IQM discards the top and bottom 25% of runs before averaging, reducing sensitivity to outlier seeds. Using these tools and stratified bootstrap confidence intervals, Agarwal et al. (2021b) find that several widely-cited algorithmic improvements on Atari 100k vanish or reverse when statistical uncertainty is accounted for.105105105They also introduce performance profiles, which plot the fraction of tasks and seeds where an algorithm achieves performance above a threshold $\tau$, as $\tau$ varies from 0 to the maximum. Performance profiles reveal the full shape of the score distribution rather than collapsing it to a single statistic.106106106The fragility extends to hyperparameters. Eimer et al. (2023) conduct a systematic study of hyperparameter sensitivity across 6 algorithms and 17 environments, finding that default hyperparameters from published papers perform competitively in the specific environments used in those papers but generalize poorly across environments. Patterson et al. (2024) synthesize these findings into an empirical design handbook, recommending at minimum 10 seeds per configuration, IQM-based comparisons, and preregistration of hyperparameter search protocols.
@@ 6.3 Value Overestimation and Spikes
Fujimoto et al. (2018) observe that Double DQN’s correction is incomplete in continuous-action settings, where the target network and online network remain correlated through shared updates. They propose Clipped Double Q-learning: compute two Q-value estimates $Q_{1},Q_{2}$ with separate networks trained on the same data, and use $y=r+\gamma\min(Q_{1}(s^{\prime},a^{\prime}),Q_{2}(s^{\prime},a^{\prime}))$ as the Bellman target. The minimum operator introduces pessimistic underestimation, which is conservative but avoids the explosive positive bias.107107107Ciosek et al. (2019) note that clipped double Q can cause excessive pessimism under high uncertainty, proposing optimistic actor-critic as a counterweight.
@@
DQN prevents outright divergence, but van Hasselt et al. (2018) find that soft divergence, defined as a temporary spike in value estimates by more than 10% followed by recovery, occurs in the majority of DQN training runs across 57 Atari games.108108108Larger networks diverge more frequently than smaller ones, counter to the usual intuition that more expressive models should generalize better. The deadly triad does not announce itself as a training failure: the value estimates may spike and recover, leaving no visible trace in the loss curve while corrupting the policy.
Kumar et al. (2021) describe a continuous manifestation of the deadly triad: implicit under-parameterization.109109109Neural networks trained with bootstrapped TD objectives progressively lose their effective rank, with fewer and fewer neurons contributing distinct directions in the representation. This rank collapse is silent (training loss continues to decrease) but the network’s ability to represent new information degrades over time, a form of capacity loss distinct from but related to plasticity loss.
@@ 6.4 Plasticity Loss and Primacy Bias
Nikishin et al. (2022) identify primacy bias as a specific cause. Because the replay buffer is filled incrementally, early transitions are oversampled relative to later ones, and the network over-fits to early environment transitions, corrupting representations throughout the remainder of training.110110110Nikishin et al. (2022) show that 100 initial “priming” steps of excessive gradient updates degrade a SAC agent’s performance for hundreds of thousands of subsequent steps. Sokar et al. (2023) measure the fraction of dormant neurons (units with near-zero activation across the replay buffer) accumulating monotonically during training. Dohare et al. (2024) find that standard deep networks lose all plasticity within a few million gradient steps in continual learning tasks.
The proposed remedies fall into three categories: periodic resets, continual backpropagation, and architectural interventions.111111111Periodic resets reinitialize the last layers of the network while retaining the replay buffer, allowing the agent to forget overfit representations without discarding experience (Nikishin et al., 2022; D’Oro et al., 2023). Continual backpropagation replaces neurons with near-zero utility at each gradient step rather than waiting for a global reset (Dohare et al., 2024). Architectural interventions—layer normalization (Lyle et al., 2025), orthogonal initialization, spectral normalization—reduce the rate at which plasticity is lost by stabilizing gradient magnitudes and preventing weight norm growth.
@@ 6.5 Implementation Dominates Algorithmic Innovation
Engstrom et al. (2020) find that PPO with the clipping mechanism disabled performs indistinguishably from the full algorithm; TRPO (Schulman et al., 2015) with the same code-level additions matches PPO.112112112The non-clipping components that suffice are: observation normalization, reward normalization, value function clipping, global gradient norm clipping, orthogonal weight initialization, and the Adam optimizer. Andrychowicz et al. (2021) identify observation normalization, orthogonal initialization, and learning rate annealing as the three choices accounting for most variance across 250,000 agents and 250 hyperparameter configurations. Huang et al. (2022) catalog 37 implementation details required to reproduce PPO on Atari;113113113These include reward clipping to [1,1][-1,1], frame stacking to 4 frames, a specific episode termination convention, and a numerically stable normalization of the advantage estimate. omitting any produces materially different results.
The most consequential implementation detail is the distinction between termination and truncation (Pardo et al., 2018). In reinforcement learning, an episode can end for two distinct reasons: termination, where the environment reaches a natural absorbing state (the pole falls in CartPole, the robot falls in locomotion tasks), and truncation, where the episode is cut short by an external time limit. At a termination, the value of the successor state is zero: $V(s_{\text{term}})=0$. At a truncation, the episode is merely paused and the successor state has non-zero value: $V(s_{\text{trunc}})\neq 0$. Treating truncated transitions as terminated substitutes zero for a non-zero bootstrap value at every time limit boundary, corrupting every Bellman update in the vicinity of episode boundaries. Pardo et al. (2018) show that this conflation degrades performance by 20–40% on standard MuJoCo benchmarks.114114114The Gymnasium API (Towers et al., 2024) enforces the distinction by returning separate terminated and truncated flags, but most pre-2022 codebases conflate them in the done flag.
@@ 6.6 Replay Buffer Pathologies and Reward Scaling
Experience replay (Lin, 1992) decouples the data collection and learning processes, allowing a single transition to be used for multiple gradient updates. The replay ratio (the number of gradient updates per environment step) governs the trade-off between sample efficiency and data staleness. Zhang and Sutton (2017) show that increasing the replay ratio beyond a modest threshold degrades performance.115115115At high replay ratios, the distribution shift between the current policy and the behavior policy that generated the stored data grows large enough to violate the off-policy assumptions of Q-learning. The relationship is non-monotone and environment-dependent, making replay buffer size a sensitive hyperparameter with no universal default.
Schaul et al. (2016) propose prioritized experience replay (PER).116116116PER samples transitions with probability proportional to the magnitude of their TD error, on the argument that high-error transitions are the most informative. Fedus et al. (2020) revisit PER on large-scale Atari experiments and find that uniform sampling from a large enough buffer matches or outperforms PER, while being simpler to implement and tune.
Reward scaling introduces a separate class of failure modes. Standard DQN (Mnih et al., 2015) clips rewards to [1,+1][-1,+1] across all environments to stabilize training. van Hasselt et al. (2016a) observe that reward clipping changes the objective: clipped rewards make all positive events equivalent regardless of magnitude, so the agent learns to maximize the frequency of positive events rather than their cumulative value. This substitution can produce policies that are locally rational under the clipped reward but qualitatively suboptimal under the true reward. van Hasselt et al. (2016a) propose PopArt as a remedy.117117117PopArt normalizes targets to have unit variance while adjusting the output layer so that the policy remains invariant to the normalization. PopArt allows consistent learning across reward scales spanning several orders of magnitude without reward clipping.
Skalse et al. (2022) formalize reward hacking and show that any non-constant proxy reward can in principle be exploited by a sufficiently capable optimizer.118118118Skalse et al. (2022) define a proxy as unhackable if increasing expected proxy return cannot decrease expected true return. Their main result states that for the set of all stochastic policies, two reward functions are unhackable only if one of them is constant. For deterministic policies and finite policy sets, non-trivial unhackable pairs exist, but the conditions are stringent.

### 23 Mar 2026 v4

@@ 5.6 Fundamental Tradeoffs
Four additional trade-offs pervade reinforcement learning. First, exploration versus exploitation: should the agent act on its current best estimate or gather information to improve future decisions? Lai and Robbins (1985) establish the fundamental lower bound: any consistent policy must incur regret at least logarithmic in the number of periods. Naive exploration ($\varepsilon$-greedy) requires samples exponential in the horizon; strategic exploration (UCB, optimism in the face of uncertainty) reduces this to polynomial (Auer et al., 2002a), formalizing the value of targeted experimentation.107107107The exploration-exploitation tradeoff is the subject of the bandits chapter, where the multi-armed bandit framework provides the sharpest analysis. In full MDPs, exploration is harder because the agent must learn not just reward distributions but also transition dynamics, compounding the information requirement.
Second, model-based versus model-free: model-based methods learn a transition model $\hat{P}(s^{\prime}|s,a)$ and plan with it (Sutton, 1990); model-free methods learn value functions or policies directly from transitions. The Dyna architecture (Sutton, 1990) bridges these by generating simulated experience from the learned model to supplement real transitions. Model-based methods are sample-efficient (each transition updates the entire model, which improves value estimates for all states) but suffer asymptotic bias if the model class is misspecified; model-free methods are asymptotically unbiased but sample-inefficient, using each transition for a single gradient step.108108108Model misspecification in RL is the analog of omitted variable bias in econometrics: if the learned model omits relevant state variables or misspecifies the functional form of transitions, the resulting policy is biased regardless of sample size. Moerland et al. (2023) provide a comprehensive survey of model-based RL, analyzing the model-bias versus sample-efficiency tradeoff across method families.
@@ 6.1 The Moving Target Problem
Fujimoto et al. (2022) formalize this observation. Let $Q^{\pi}$ denote the true value function for policy $\pi$, let $\Delta(s,a)=Q(s,a)-Q^{\pi}(s,a)$ denote the value error, and let $\varepsilon(s,a)=Q(s,a)-(r+\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[Q(s^{\prime},a^{\prime})])$ denote the Bellman error. Substituting the definition of $Q^{\pi}$ yields the key identity: $\varepsilon(s,a)=\Delta(s,a)-\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[\Delta(s^{\prime},a^{\prime})]$. The Bellman error is a difference of value errors at consecutive states, not the value error itself. If the errors $\Delta(s,a)$ and $\Delta(s^{\prime},a^{\prime})$ are correlated across time—the network is wrong in the same direction at successive state-action pairs—they cancel in the difference and the Bellman error is small regardless of how large the individual errors are.109109109In the extreme case, shifting all Q-values by the constant $c/(1-\gamma)$ leaves the Bellman error unchanged at zero while increasing value error by $c/(1-\gamma)$.
The second failure mode is specific to finite datasets: Fujimoto et al. (2022, Corollary 1) show that over an incomplete dataset, zero Bellman error is consistent with arbitrarily large value error, because the network can fit unobserved successor pairs to whatever values make the observed residuals vanish.110110110The Bellman equation uniquely identifies $Q^{\pi}$ when enforced over the entire MDP, but over an incomplete dataset it admits infinitely many solutions. Whenever a transition $(s^{\prime},a^{\prime})$ that is reachable from the dataset is not itself in the dataset, the network is free to set $Q(s^{\prime},a^{\prime})$ at unobserved pairs to whatever value makes the residual vanish on the observed ones, unconstrained by any loss. A network can thus reach near-zero training loss while the value function remains arbitrarily inaccurate over the full state space.
The dual failure mode appears in policy gradient methods: Ilyas et al. (2020) find that even when PPO’s surrogate objective improves monotonically, episode return can plateau or decline, because the surrogate gradient is poorly aligned with the gradient of the true return.111111111Ilyas et al. (2020) examine the surrogate objective used in Proximal Policy Optimization (Schulman et al., 2017). The PPO clipping mechanism is designed to keep policy updates within a trust region by bounding the probability ratio $\pi_{\theta}(a|s)/\pi_{\theta_{\text{old}}}(a|s)$. The gradient of the surrogate is poorly aligned with the gradient of the true return, particularly in later training phases where the policy has diverged from the behavior policy used to collect the replay data. The loss metric that practitioners monitor throughout training is measuring something other than what they care about.
@@ 6.2 The Reproducibility Crisis and Sensitivity to Random Seeds
Henderson et al. (2018) trained five leading policy gradient algorithms (PPO, TRPO, DDPG, TD3, SAC) on six MuJoCo benchmark environments, holding all hyperparameters fixed and varying only the random seed. The resulting learning curves from different seeds were non-overlapping: a seed that performed well under one algorithm performed comparably to a different algorithm’s best seeds, making cross-algorithm comparison unreliable.112112112Differences as large as 2,000 points in final episode return arose from seed variation alone. Agarwal et al. (2021b) quantify the damage: comparing point estimates from 5 runs per task on Atari 100k yields Type I error exceeding 50%, meaning a random noise injection appears beneficial in half of all comparisons.
Agarwal et al. (2021b) propose the interquartile mean (IQM) as a replacement for mean and median when comparing algorithms. The IQM discards the top and bottom 25% of runs before averaging, reducing sensitivity to outlier seeds. Using these tools and stratified bootstrap confidence intervals, Agarwal et al. (2021b) find that several widely-cited algorithmic improvements on Atari 100k vanish or reverse when statistical uncertainty is accounted for.113113113They also introduce performance profiles, which plot the fraction of tasks and seeds where an algorithm achieves performance above a threshold $\tau$, as $\tau$ varies from 0 to the maximum. Performance profiles reveal the full shape of the score distribution rather than collapsing it to a single statistic.114114114The fragility extends to hyperparameters. Eimer et al. (2023) conduct a systematic study of hyperparameter sensitivity across 6 algorithms and 17 environments, finding that default hyperparameters from published papers perform competitively in the specific environments used in those papers but generalize poorly across environments. Patterson et al. (2024) synthesize these findings into an empirical design handbook, recommending at minimum 10 seeds per configuration, IQM-based comparisons, and preregistration of hyperparameter search protocols.
@@ 6.3 Value Overestimation and Spikes
Fujimoto et al. (2018) observe that Double DQN’s correction is incomplete in continuous-action settings, where the target network and online network remain correlated through shared updates. They propose Clipped Double Q-learning: compute two Q-value estimates $Q_{1},Q_{2}$ with separate networks trained on the same data, and use $y=r+\gamma\min(Q_{1}(s^{\prime},a^{\prime}),Q_{2}(s^{\prime},a^{\prime}))$ as the Bellman target. The minimum operator introduces pessimistic underestimation, which is conservative but avoids the explosive positive bias.115115115Ciosek et al. (2019) note that clipped double Q can cause excessive pessimism under high uncertainty, proposing optimistic actor-critic as a counterweight.
@@
DQN prevents outright divergence, but van Hasselt et al. (2018) find that soft divergence, defined as a temporary spike in value estimates by more than 10% followed by recovery, occurs in the majority of DQN training runs across 57 Atari games.116116116Larger networks diverge more frequently than smaller ones, counter to the usual intuition that more expressive models should generalize better. The deadly triad does not announce itself as a training failure: the value estimates may spike and recover, leaving no visible trace in the loss curve while corrupting the policy.
Kumar et al. (2021) describe a continuous manifestation of the deadly triad: implicit under-parameterization.117117117Neural networks trained with bootstrapped TD objectives progressively lose their effective rank, with fewer and fewer neurons contributing distinct directions in the representation. This rank collapse is silent (training loss continues to decrease) but the network’s ability to represent new information degrades over time, a form of capacity loss distinct from but related to plasticity loss.
@@ 6.4 Plasticity Loss and Primacy Bias
Nikishin et al. (2022) identify primacy bias as a specific cause. Because the replay buffer is filled incrementally, early transitions are oversampled relative to later ones, and the network over-fits to early environment transitions, corrupting representations throughout the remainder of training.118118118Nikishin et al. (2022) show that 100 initial “priming” steps of excessive gradient updates degrade a SAC agent’s performance for hundreds of thousands of subsequent steps. Sokar et al. (2023) measure the fraction of dormant neurons (units with near-zero activation across the replay buffer) accumulating monotonically during training. Dohare et al. (2024) find that standard deep networks lose all plasticity within a few million gradient steps in continual learning tasks.
The proposed remedies fall into three categories: periodic resets, continual backpropagation, and architectural interventions.119119119Periodic resets reinitialize the last layers of the network while retaining the replay buffer, allowing the agent to forget overfit representations without discarding experience (Nikishin et al., 2022; D’Oro et al., 2023). Continual backpropagation replaces neurons with near-zero utility at each gradient step rather than waiting for a global reset (Dohare et al., 2024). Architectural interventions—layer normalization (Lyle et al., 2025), orthogonal initialization, spectral normalization—reduce the rate at which plasticity is lost by stabilizing gradient magnitudes and preventing weight norm growth.
@@ 6.5 Implementation Dominates Algorithmic Innovation
Engstrom et al. (2020) find that PPO with the clipping mechanism disabled performs indistinguishably from the full algorithm; TRPO (Schulman et al., 2015) with the same code-level additions matches PPO.120120120The non-clipping components that suffice are: observation normalization, reward normalization, value function clipping, global gradient norm clipping, orthogonal weight initialization, and the Adam optimizer. Andrychowicz et al. (2021) identify observation normalization, orthogonal initialization, and learning rate annealing as the three choices accounting for most variance across 250,000 agents and 250 hyperparameter configurations. Huang et al. (2022) catalog 37 implementation details required to reproduce PPO on Atari;121121121These include reward clipping to [1,1][-1,1], frame stacking to 4 frames, a specific episode termination convention, and a numerically stable normalization of the advantage estimate. omitting any produces materially different results.
The most consequential implementation detail is the distinction between termination and truncation (Pardo et al., 2018). In reinforcement learning, an episode can end for two distinct reasons: termination, where the environment reaches a natural absorbing state (the pole falls in CartPole, the robot falls in locomotion tasks), and truncation, where the episode is cut short by an external time limit. At a termination, the value of the successor state is zero: $V(s_{\text{term}})=0$. At a truncation, the episode is merely paused and the successor state has non-zero value: $V(s_{\text{trunc}})\neq 0$. Treating truncated transitions as terminated substitutes zero for a non-zero bootstrap value at every time limit boundary, corrupting every Bellman update in the vicinity of episode boundaries. Pardo et al. (2018) show that this conflation degrades performance by 20–40% on standard MuJoCo benchmarks.122122122The Gymnasium API (Towers et al., 2024) enforces the distinction by returning separate terminated and truncated flags, but most pre-2022 codebases conflate them in the done flag.
@@ 6.6 Replay Buffer Pathologies and Reward Scaling
Experience replay (Lin, 1992) decouples the data collection and learning processes, allowing a single transition to be used for multiple gradient updates. The replay ratio (the number of gradient updates per environment step) governs the trade-off between sample efficiency and data staleness. Zhang and Sutton (2017) show that increasing the replay ratio beyond a modest threshold degrades performance.123123123At high replay ratios, the distribution shift between the current policy and the behavior policy that generated the stored data grows large enough to violate the off-policy assumptions of Q-learning. The relationship is non-monotone and environment-dependent, making replay buffer size a sensitive hyperparameter with no universal default.
Schaul et al. (2016) propose prioritized experience replay (PER).124124124PER samples transitions with probability proportional to the magnitude of their TD error, on the argument that high-error transitions are the most informative. Fedus et al. (2020) revisit PER on large-scale Atari experiments and find that uniform sampling from a large enough buffer matches or outperforms PER, while being simpler to implement and tune.
Reward scaling introduces a separate class of failure modes. Standard DQN (Mnih et al., 2015) clips rewards to [1,+1][-1,+1] across all environments to stabilize training. van Hasselt et al. (2016a) observe that reward clipping changes the objective: clipped rewards make all positive events equivalent regardless of magnitude, so the agent learns to maximize the frequency of positive events rather than their cumulative value. This substitution can produce policies that are locally rational under the clipped reward but qualitatively suboptimal under the true reward. van Hasselt et al. (2016a) propose PopArt as a remedy.125125125PopArt normalizes targets to have unit variance while adjusting the output layer so that the policy remains invariant to the normalization. PopArt allows consistent learning across reward scales spanning several orders of magnitude without reward clipping.
Skalse et al. (2022) formalize reward hacking and show that any non-constant proxy reward can in principle be exploited by a sufficiently capable optimizer.126126126Skalse et al. (2022) define a proxy as unhackable if increasing expected proxy return cannot decrease expected true return. Their main result states that for the set of all stochastic policies, two reward functions are unhackable only if one of them is constant. For deterministic policies and finite policy sets, non-trivial unhackable pairs exist, but the conditions are stringent.